ALLEN BOUGHT 20 STAMPS IN 37 CENT AND 20 CENT DENOMONATIONS. IF THE TOTAL COST OF THE STAMPS WAS $7.06, HOW MANY 37 CENT STAMPS DID ALLEN BUY
Answer by checkley71(8403)

X+Y=20 OR X=20-Y
.37X+.20Y=7.06
NOW REPLACE X WITH(20-Y)AND SOLVE FOR Y IN THE SECOND EQUATION
.37(20-Y)+20Y=7.06
7.40-37Y+20Y=7.06
-17Y=7.40-7.06
-17Y=-.34
Y=-34/-17
Y=2 $.20 STAMPS.
20-2=18 $.37 STAMPS.
PROOF
.37*18+2*.20=7.06
6.66+.40=7.06
7.06=7.06